# This information aligns with the guidelines available at:

### v1

@@ Abstract
Previous studies demonstrate the impressive performance of residual neural networks (ResNet) in speaker verification. The ResNet models treat the time and frequency dimensions equally. They follow the default stride configuration designed for image recognition, where the horizontal and vertical axes exhibit similarities. This approach ignores the fact that time and frequency are asymmetric in speech representation. In this paper, we address this issue and look for optimal stride configurations specifically tailored for speaker verification. We represent the stride space on a trellis diagram, and conduct a systematic study on the impact of temporal and frequency resolutions on the performance and further identify two optimal points, namely Golden Gemini, which serves as a guiding principle for designing 2D ResNet-based speaker verification models. By following the principle, a state-of-the-art ResNet baseline model gains a significant performance improvement on VoxCeleb, SITW, and CNCeleb datasets with 7.70%/11.76% average EER/minDCF reductions, respectively, across different network depths (ResNet18, 34, 50, and 101), while reducing the number of parameters by 16.5% and FLOPs by 4.1%. We refer to it as Gemini ResNet. Further investigation reveals the efficacy of the proposed Golden Gemini operating points across various training conditions and architectures. Furthermore, we present a new benchmark, namely the Gemini DF-ResNet, using a cutting-edge model. Codes and pre-trained models will be available at https://github.com/wenet-e2e/wespeaker.
@@ I Introduction
The ResNet architecture was initially designed for image recognition [50] where the horizontal and vertical dimensions of images have similar implications [50, 69, 70] and are often uniform in size, typically $N\times N$ pixels with commonly used values such as 224 and 384. Consequently, it is intuitive to treat these two dimensions equally with the default equal-stride configuration in ResNet [50]. However, when dealing with speech representations, the time and frequency axes of speech spectrograms possess distinct implications [71] and often vary in size (e.g., 80 $\times$ 301 [40]). Therefore, the techniques that work for image recognition may not be suitable for ASV, thus necessitating appropriate modifications. Despite these notable differences in feature properties between images and speech signals, existing ASV systems based on the ResNet models [19, 20, 31, 18, 23, 21, 51, 52, 53, 54, 22, 28, 30, 24, 25, 26, 27, 29] continue to treat the frequency and temporal resolutions equally by adopting the default stride configuration as the original ResNet [50]. Doubts arise regarding the adequacy of this equal-stride configuration for ASV.
The preservation of temporal resolution in various existing ASV methods [72, 39, 32, 33, 34, 35, 36, 31, 42, 40, 41, 38, 37, 73, 74, 47] has led to the hypothesis that ASV may be more sensitive to temporal resolution than frequency resolution. TDNN-based models [39, 32, 33, 34, 35, 36, 31, 40, 41, 42, 47] preserve the temporal resolution across the stacked layers. Similarly, recurrent networks, such as the long short-term memory (LSTM), preserve the number of frames [73, 74]. Recent studies  [38, 37, 47] adopt the Transformer architecture as the encoder, ensuring the preservation of the temporal resolution across stacked Transformer blocks. Should temporal resolution prove to be of greater significance, the equal-stride configuration may not be optimal since it diminishes the temporal resolution. The current understanding of the joint impact of temporal and frequency resolutions on the performance of ResNet-based ASV remains limited. Consequently, this motivates us to explore the relative importance of temporal and frequency resolution in the feature representation process of ASV. Building upon this investigation, we identify the optimal stride configurations that account for the inherent characteristics of speech signals to better align with the requirements of ASV, leading to improved performance. We also conduct a meticulous analysis of the trade-offs between performance and model complexity to ensure both efficacy and efficiency. The major contributions of this work are summarized as follows:
@@ II-C Stride and Resolution
As illustrated in Fig. 1 (a), a TDNN network implemented with dilated 1D CNN layers [32] treats the input as 1D features, while considering the frequency dimension as channels [32, 39, 40]. Unlike TDNNs, a 2D CNN considers the input feature as a 3-dimensional tensor $C\times F\times T$, where C, F, and T represent the channel, frequency, and time dimensions, respectively [24]. By employing multiple 2D CNN layers, the number of channels increases, while the frequency and temporal resolutions decrease by downsampling operations to reduce computational complexity [50]. The output dimension of the downsampling operation is mainly controlled by the stride. Figure 1 (b) and (c) illustrate that by adjusting the stride in each dimension, the temporal and frequency resolutions can be controlled independently. For instance, setting the stride to 2 on the time dimension and 1 for the frequency dimension roughly halves the time resolution while keeping the frequency resolution the same.
@@ IV-B Training Strategy
Training strategy 2: Wespeaker Toolkit [^3] https://github.com/wenet-e2e/wespeaker [21] is used. This training strategy follows that in  [24] for the purpose of re-implementing DF-ResNet [24] and is only applied to re-implemented DF-ResNet and Gemini DF-ResNet reported in Section V-E. Specifically, the loss function is an AAM-softmax [16] with a margin of 0.2 and a scale of 32. The total number of training epochs is set to 165. The AdamW [91] optimizer with 0.05 weight decay is used. The base learning rate ($l_{\mathrm{base}}$) decreases from $1.25\times 10^{-4}$ to $1\times 10^{-6}$ with the exponential scheduler as the learning rate regulator. The learning rate ($l$) for training is adjusted according to the batch size ($b$) and formulated as $l=l_{\mathrm{base}}\times b/64$. In addition, we use an exponential warm-up [50] of the learning rate in the first six epochs. All the samples are cut into 200-frame segments with the augmentations of reverberation, noise, and speed perturbation [88] during training. The as-norm [92] is applied to normalize the scores.
@@ V-B Finding the Sweet Spots on the Trellis Diagram
Downsampling Stride Config. Vox1-O Factors [Time] FLOPs (Giga) EER Index of Stride Config. ($\alpha_{\mathrm{5}}$, $\beta_{\mathrm{5}}$) [Frequency] Params (Million) 2s/3s minDCF [2,2,2,2,2] 2.744 ORI (32,32) [2,2,2,2,2] 21.41 1.25/1.82 0.291 [1,1,2,2,2] 1.101 MOD (8,8) [1,1,2,2,2] 6.63 4.63/6.88 0.128 [1,1,1,1,1] 1.250 T05 (1,32) [2,2,2,2,2] 5.72 4.49/6.72 0.131 [2,2,2,2,2] 2.526 F50 (32,1) [1,1,1,1,1] 15.81 4.44/6.43 0.241 [1,1,1,1,2] 1.303 T15 (2,32) [2,2,2,2,2] 5.72 3.50/5.24 0.161 [2,2,2,2,2] 2.505 F51 (32,2) [1,1,1,1,2] 10.57 3.52/5.11 0.243 [1,1,1,2,2] 1.218 T25 (4,32) [2,2,2,2,2] 5.72 2.16/3.23 0.133 [2,2,2,2,2] 2.228 F52 (32,4) [1,1,1,2,2] 7.95 2.17/3.16 0.219 [1,1,1,1,2] 1.058 T14 (2,16) [1,2,2,2,2] 5.98 6.68/9.99 0.092 [1,2,2,2,2] 1.882 F41 (16,2) [1,1,1,1,2] 10.57 6.87/10.08 0.150 [1,1,1,2,2] 1.111 T24 (4,16) [1,2,2,2,2] 5.98 4.15/6.20 0.104 [1,2,2,2,2] 1.563 F42 (16,4) [1,1,1,2,2] 7.95 4.24/6.24 0.135 [1,1,2,2,2] 1.260 T34 (8,16) [1,2,2,2,2] 5.98 2.32/3.46 0.128 [1,2,2,2,2] 1.691 F43 (16,8) [1,1,2,2,2] 6.64 2.35/3.47 0.180 [1,1,1,2,2] 1.101 T23 (4,8) [1,1,2,2,2] 6.63 8.27/12.37 0.095 [1,1,2,2,2] 1.223 F32 (8,4) [1,1,1,2,2] 7.95 8.35/12.41 0.105 [1,1,1,1,1] 1.276 T04 (1,16) [1,2,2,2,2] 5.98 8.32/12.45 0.117 [1,1,1,1,2] 1.127 T13 (2,8) [1,1,2,2,2] 6.63 13.33/19.95 0.107
@@
Observation 2 – The performance of models with endpoints located on the boundary will significantly deteriorate. As shown in Table V-B, models indexed as T05, T15, T25, F52, F51, and F50 exhibit notable performance degradation compared to neighboring models on the trellis diagram. Unlike a TDNN that utilizes large channel numbers (e.g., 512, 1024, or 2048) [32, 40, 39], ResNet employs a smaller channel number (such as 32 or 64) at early stages for low-dimensional information representations [19, 24, 25, 21]. This aligns with the design principle discussed in Section II-B that emphasizes the importance of depth over width. Consequently, when the temporal or frequency resolution is rapidly compressed, and constrained by a limited number of filters, it leads to the loss of information in that specific dimension. This results in a notable degradation of performance. Therefore, when designing a narrow ResNet with a smaller width, it is advisable to avoid the stride configurations located on the boundary.
@@ V-D Evaluation on Compatibility
This subsection aims to confirm the consistently superior performance of the proposed Golden-Gemini stride configuration across various conditions and its compatibility with different techniques. The experiments are conducted to compare the models using the Golden Gemini T14 stride configuration and the default equal-stride configuration under the following conditions:
@@ V-E New SOTA Benchmark
DF-ResNet [24, 25] is a series of powerful SOTA models introduced in Section II-B. We first re-implement a small venison, namely DF-ResNet59. The results reported in Table 6 demonstrate that the re-implemented DF-ResNet model slightly outperforms the one reported in [24, 25]. Notably, we do not apply SpecAugment [87] which is used in [24, 25]. SpecAugment has been proven effective in automatic speech recognition (ASR) [87]. However, it can have adverse effects on the fundamental frequency of the audio, which is a critical characteristic for speaker discrimination [96]. Prior work [21] shows that combining SpecAugment with other augmentation methods in ASV can pose compatibility challenges. Our experiments demonstrate a similar trend.
Similar to other ResNet models, DF-ResNet initially adopts the default equal-stride configuration, treating temporal and frequency dimensions equally. By replacing the equal-stride configuration with our proposed Golden-Gemini stride configuration T14c, we achieve significant performance improvements and reduce model complexity, as demonstrated in Table VIII and 6. Table 6 also includes a comparison of the proposed Gemini DF-ResNet models with other SOTA systems. The comparison unequivocally demonstrates that the proposed Gemini DF-ResNet achieves the best performance among all the systems to the best of our knowledge. This outcome significantly underscores the remarkable capabilities of the Golden-Gemini stride configuration and support Golden-Gemini Hypothesis, which emphasizes the critical significance of temporal resolution in attaining superior results in ASV.
@@
TABLE IX: Performance in EER(%) and minDCF of the proposed Golden Gemini DF-ResNet and SOTA systems on VoxCeleb1 test sets. Models with our proposed Golden-Gemini stride configuration are highlighted in grey  666 Codes and pre-trained models of the proposed Gemini DF-ResNet would be available at https://github.com/wenet-e2e/wespeaker..
@@ VI Conclusion
We investigate efficient stride configurations for speaker verification. Through a strategic search on a trellis diagram, we analyze the impact of temporal and frequency resolution on the ASV performance. Experimental results on the VoxCeleb, SITW, and CNCeleb test sets highlight the significance of the temporal resolution. This leads us to identify two points, named Golden Gemini, representing two series of optimal stride configurations for ASV. We also present a set of guiding principles that comprehensively describe the Golden Gemini for designing 2D ResNet for ASV. Further experiments demonstrate the consistent superiority and excellent compatibility of the proposed Golden Gemini with various structures across different conditions. Moreover, our approach is simple yet effective and can be easily applied to any 2D ResNet architecture style, offering improved performance while reducing model complexity. Based on the Golden-Gemini guiding principles, we introduce a powerful benchmark for ASV, namely the Gemini DF-ResNet. These findings indicate the promising value of our method in real-world applications and its great potential in related tasks, such as speaker diarization, speaker extraction, and speech anti-spoofing.

### v3

# Golden Gemini is All You Need: Finding the Sweet Spots for Speaker Verification

@@ Abstract
The residual neural networks (ResNet) demonstrate the impressive performance in automatic speaker verification (ASV). They treat the time and frequency dimensions equally, following the default stride configuration designed for image recognition, where the horizontal and vertical axes exhibit similarities. This approach ignores the fact that time and frequency are asymmetric in speech representation. We address this issue and postulate Golden-Gemini Hypothesis, which posits the prioritization of temporal resolution over frequency resolution for ASV. The hypothesis is verified by conducting a systematic study on the impact of temporal and frequency resolutions on the performance, using a trellis diagram to represent the stride space. We further identify two optimal points, namely Golden Gemini, which serves as a guiding principle for designing 2D ResNet-based ASV models. By following the principle, a state-of-the-art ResNet baseline model gains a significant performance improvement on VoxCeleb, SITW, and CNCeleb datasets with 7.70%/11.76% average EER/minDCF reductions, respectively, across different network depths (ResNet18, 34, 50, and 101), while reducing the number of parameters by 16.5% and FLOPs by 4.1%. We refer to it as Gemini ResNet. Further investigation reveals the efficacy of the proposed Golden Gemini operating points across various training conditions and architectures. Furthermore, we present a new benchmark, namely the Gemini DF-ResNet, using a cutting-edge model. Codes and pre-trained models are available at https://github.com/Tianchi-Liu9/Golden-Gemini-for-Speaker-Verification.
@@ I Introduction
The ResNet architecture was initially designed for image recognition [50] where the horizontal and vertical dimensions of images have similar implications [69, 70] and are often uniform in size, typically $N\times N$ pixels with commonly used values such as 224 and 384. Consequently, it is intuitive to treat these two dimensions equally with the default equal-stride configuration in ResNet [50]. However, when dealing with speech representations, the time and frequency axes of speech spectrograms possess distinct implications [71] and often vary in size (e.g., 80 $\times$ 301 [40]). Therefore, the techniques that work for image recognition may not be suitable for ASV, thus necessitating appropriate modifications. Despite these notable differences in feature properties between images and speech signals, existing ASV systems based on the ResNet models [19, 20, 31, 18, 23, 21, 51, 52, 53, 54, 22, 28, 30, 24, 25, 26, 27, 29] continue to treat the frequency and temporal resolutions equally by adopting the default stride configuration as the original ResNet. Doubts arise regarding the adequacy of this equal-stride configuration for ASV.
The preservation of temporal resolution in various existing ASV methods [72, 39, 32, 33, 34, 35, 36, 31, 42, 40, 41, 38, 37, 73, 74, 47] has led to the hypothesis that ASV may be more sensitive to temporal resolution than frequency resolution. TDNN-based models [39, 32, 33, 34, 35, 36, 31, 40, 41, 42, 47] preserve the temporal resolution across the stacked layers. Similarly, recurrent networks, such as the long short-term memory (LSTM), preserve the number of frames [73, 74]. Recent studies  [38, 37, 47] adopt the Transformer architecture as the encoder, ensuring the preservation of the temporal resolution across stacked Transformer blocks. Should temporal resolution prove to be of greater significance, the equal-stride configuration may not be optimal since it diminishes the temporal resolution. The current understanding of the impact of temporal and frequency resolutions on the performance of ResNet-based ASV models remains limited, leaving a research gap to be filled. Consequently, this motivates us to explore the relative importance of temporal and frequency resolution in the feature representation process of ASV. Building upon this investigation, we identify the optimal stride configurations that account for the inherent characteristics of speech signals to better align with the requirements of ASV, leading to improved performance. We also conduct a meticulous analysis of the trade-offs between performance and model complexity to ensure both efficacy and efficiency. The major contributions of this work are summarized as follows:
  • We postulate Golden-Gemini Hypothesis, which posits that the preservation of temporal resolution is to be prioritized over frequency resolution for the optimal extraction of speaker characteristics.
@@ II-C Stride and Resolution
As illustrated in Fig. 1 (a), a TDNN network implemented with dilated 1D CNN layers [32] treats the input as 1D features, while considering the frequency dimension as channels. TDNN-based models are not included in this study due to the absence of the frequency dimension, and existing TDNN models generally maintain time resolution [39, 32, 33, 34, 35, 36, 31, 40, 41, 42, 47]. Unlike TDNNs, a 2D CNN considers the input feature as a 3-dimensional tensor $C\times F\times T$, where C, F, and T represent the channel, frequency, and time dimensions, respectively [24]. By employing multiple 2D CNN layers, the number of channels increases, while the frequency and temporal resolutions decrease by downsampling operations to reduce computational complexity [50]. The output dimension of the downsampling operation is mainly controlled by the stride. Figure 1 (b) and (c) illustrate that by adjusting the stride in each dimension, the temporal and frequency resolutions can be controlled independently. For instance, setting the stride to 2 on the time dimension and 1 for the frequency dimension roughly halves the time resolution while keeping the frequency resolution the same.
@@ IV-B Training Strategy
Training strategy 2: Wespeaker Toolkit [^3] https://github.com/wenet-e2e/wespeaker [21] is used. This training strategy follows that in  [24] for the purpose of re-implementing DF-ResNet [24] and is only applied to re-implemented DF-ResNet and Gemini DF-ResNet reported in Section V-E. Specifically, the loss function is an AAM-softmax [16] with a margin of 0.2 and a scale of 32. The total number of training epochs is set to 165. The AdamW [91] optimizer with 0.05 weight decay is used. The base learning rate ($l_{\mathrm{base}}$) decreases from $1.25\times 10^{-4}$ to $1\times 10^{-6}$ with the exponential scheduler as the learning rate regulator. The learning rate ($l$) for training is adjusted according to the batch size ($b$) and formulated as $l=l_{\mathrm{base}}\times b/64$. All the samples are cut into 200-frame segments with the augmentations of reverberation, noise, and speed perturbation [88] during training. The as-norm [92] is applied to normalize the scores.
@@ V-B Finding the Sweet Spots on the Trellis Diagram
Downsampling Stride Config. Vox1-O
Factors [Time] FLOPs (Giga) EER
[2,2,2,2,2] 2.744
ORI (32,32) [2,2,2,2,2] 21.41 1.25/1.82 0.291
[1,1,2,2,2] 1.101
MOD (8,8) [1,1,2,2,2] 6.63 4.63/6.88 0.128
[1,1,1,1,1] 1.250
T05 (1,32) [2,2,2,2,2] 5.72 4.49/6.72 0.131
[2,2,2,2,2] 2.526
F50 (32,1) [1,1,1,1,1] 15.81 4.44/6.43 0.241
[1,1,1,1,2] 1.303
T15 (2,32) [2,2,2,2,2] 5.72 3.50/5.24 0.161
[2,2,2,2,2] 2.505
F51 (32,2) [1,1,1,1,2] 10.57 3.52/5.11 0.243
[1,1,1,2,2] 1.218
T25 (4,32) [2,2,2,2,2] 5.72 2.16/3.23 0.133
[2,2,2,2,2] 2.228
F52 (32,4) [1,1,1,2,2] 7.95 2.17/3.16 0.219
[1,1,1,1,2] 1.058
T14 (2,16) [1,2,2,2,2] 5.98 6.68/9.99 0.092
[1,2,2,2,2] 1.882
F41 (16,2) [1,1,1,1,2] 10.57 6.87/10.08 0.150
[1,1,1,2,2] 1.111
T24 (4,16) [1,2,2,2,2] 5.98 4.15/6.20 0.104
[1,2,2,2,2] 1.563
F42 (16,4) [1,1,1,2,2] 7.95 4.24/6.24 0.135
[1,1,2,2,2] 1.260
T34 (8,16) [1,2,2,2,2] 5.98 2.32/3.46 0.128
[1,2,2,2,2] 1.691
F43 (16,8) [1,1,2,2,2] 6.64 2.35/3.47 0.180
[1,1,1,2,2] 1.101
T23 (4,8) [1,1,2,2,2] 6.63 8.27/12.37 0.095
[1,1,2,2,2] 1.223
F32 (8,4) [1,1,1,2,2] 7.95 8.35/12.41 0.105
[1,1,1,1,1] 1.276
T04 (1,16) [1,2,2,2,2] 5.98 8.32/12.45 0.117
[1,1,1,1,2] 1.127
T13 (2,8) [1,1,2,2,2] 6.63 13.33/19.95 0.107
@@
Observation 2 – The performance of models with endpoints located on the boundary will significantly deteriorate. As shown in Table IV, models indexed as T05, T15, T25, F52, F51, and F50 exhibit notable performance degradation compared to neighboring models on the trellis diagram. Unlike a TDNN that utilizes large channel numbers (e.g., 512, 1024, or 2048) [32, 40, 39], ResNet employs a smaller channel number (such as 32 or 64) at early stages for low-dimensional information representations [19, 24, 25, 21]. This aligns with the design principle discussed in Section II-B that emphasizes the importance of depth over width. Consequently, when the temporal or frequency resolution is rapidly compressed, and constrained by a limited number of filters, it leads to the loss of information in that specific dimension. This results in a notable degradation of performance. Therefore, when designing a narrow ResNet with a smaller width, it is advisable to avoid the stride configurations located on the boundary.
@@ V-D Evaluation on Compatibility
The changes in time and frequency resolutions occur once per stage in both ResNet and its variant networks, such as DF-ResNet [24], Res2Net [76], and SD-ResNet [70]. Given that the Golden Gemini is concluded from investigating the significance of time and frequency resolution for speaker verification, it is expected to apply to all ResNet series networks that still adhere to the original ResNet’s five-stage structural design. This subsection aims to confirm the consistently superior performance of the proposed Golden-Gemini stride configuration across various conditions and its compatibility with different techniques. We exemplify with the Golden Gemini T14c stride configuration, conducting experiments to compare the models using the Golden Gemini T14c and the default equal-stride configuration under the following conditions:
@@ V-E New SOTA Benchmark
DF-ResNet [24, 25] is a series of powerful SOTA models introduced in Section II-B. We first re-implement a small venison, namely DF-ResNet59. The results reported in Table IX demonstrate that the re-implemented DF-ResNet model slightly outperforms the one reported in [24, 25]. Notably, we do not apply SpecAugment [87] which is used in [24, 25]. SpecAugment has been proven effective in automatic speech recognition (ASR) [87]. However, it can have adverse effects on the fundamental frequency of the audio, which is a critical characteristic for speaker discrimination [96]. Prior work [21] shows that combining SpecAugment with other augmentation methods in ASV can pose compatibility challenges. Our experiments demonstrate a similar trend.
@@
TABLE IX: Performance in EER(%) and minDCF of the proposed Golden Gemini DF-ResNet and SOTA systems on VoxCeleb1 test sets. Models with our proposed Golden-Gemini stride configuration are highlighted in grey  666 Pre-trained models and codes of the proposed Gemini DF-ResNet are available at https://github.com/Tianchi-Liu9/Golden-Gemini-for-Speaker-Verification and https://github.com/wenet-e2e/wespeaker..
@@
Similar to other ResNet models, DF-ResNet adopts the default equal-stride configuration, treating temporal and frequency dimensions equally. By replacing the stride configuration with our proposed Golden-Gemini T14c, we see a notable 4.9% average performance boost and a 7.6% reduction in model size, as detailed in Table IX. Also, Table VIII shows a 4.5% decrease in FLOPs. It’s important to note that DF-ResNet, our chosen baseline, achieves SOTA performance with a relatively small model size, emphasizing its meticulous design and efficiency. In this context, achieving further performance gains becomes challenging given the already very low EER and minDCF. Further analysis indicates that, as the model size increases from the smallest to largest, relative performance improvements decrease from 7.4% to 5.8%, and then to 1.7%. This trend aligns with the inherent difficulty of achieving significant performance improvements over a robust baseline and low EER/minDCF. Nevertheless, our proposed Golden-Gemini stride configuration still brings improvements, securing the best performance among all systems. This outcome supports the remarkable capabilities of the Golden-Gemini stride configuration and Golden-Gemini Hypothesis, which emphasizes the critical significance of temporal resolution in attaining superior results in ASV.
@@ VI Conclusion
We investigate efficient stride configurations for speaker verification. Through a strategic search on a trellis diagram, we analyze the impact of temporal and frequency resolution on the ASV performance. Experimental results on the VoxCeleb, SITW, and CNCeleb test sets highlight the significance of the temporal resolution. This leads us to identify two points, named Golden Gemini, representing two series of optimal stride configurations for ASV. We also present a set of guiding principles that comprehensively describe the Golden Gemini for designing 2D ResNet for ASV. Further experiments demonstrate the consistent superiority and excellent compatibility of the proposed Golden Gemini with various structures across different conditions. Moreover, our approach is simple yet effective and can be easily applied to any 2D ResNet architecture style, offering improved performance while reducing model complexity. Based on the Golden-Gemini guiding principles, we introduce a powerful benchmark for ASV, namely the Gemini DF-ResNet. These findings indicate the promising value of our method in real-world applications. Additionally, the significance of time and frequency resolutions may extend beyond speaker verification, holding great potential for related tasks such as speaker diarization, speaker extraction, emotion recognition, and speech anti-spoofing.